OpenStudy (anonymous):

can someone help please Write the equation of a line in point-slope form that is parallel to the line y = 3x +6 and passes through the point (-5, 8)

OpenStudy (anonymous):

@CalebBeavers

OpenStudy (anonymous):

@jim_thompson5910 can u help please?

OpenStudy (anonymous):

sfsdfs

OpenStudy (anonymous):

@pipucho wat?

OpenStudy (anonymous):

Whats the slope of y=3x+6?

OpenStudy (anonymous):

3x?

OpenStudy (anonymous):

Just the number in front of x, so 3. Parallel lines have the same slope. so now use y-y1=m(x-x1) with slope 3 and the point (-5,8)

OpenStudy (anonymous):

y-8=3(x-5)?

OpenStudy (anonymous):

y-8=3(x+5) 5 changed to positve